# CodeUltraFeedback: An LLM-as-a-Judge Dataset for Aligning Large Language Models to Coding Preferences

### v1

@@ III Evaluating LLMs on CodeUltraFeedback
While CodeUltraFeedback has the potential to facilitate diverse downstream applications, including LLM alignment through RLAIF and DPO, we begin our experiments by analyzing the dataset itself. Specifically, the goal is to gauge GPT-3.5-Turbo’s judging capabilities and the innate ability of LLMs to align to coding preferences without any tailored optimization involved.
III-A LLMs Scores Exploration
We start by exploring the ratings generated by GPT-3.5-Turbo in the LLM-as-a-Judge annotation phase. In Table III, we present a detailed analysis of the average scores across coding preferences for each LLM, offering initial insight into their baseline capabilities.
First, we observe that GPT-3.5-Turbo and GPT-4-Turbo score the highest across all preferences and on average. This result is expected considering both models have undergone extensive instruction and RLHF tunings, which help align the LLMs more closely with human preferences [33].
At first glance, the scores might suggest marginal score discrepancies given the 1–5 scoring range. However, these variations are substantial and carry statistical significance. To quantify this significance, we performed pairwise Welch’s $t$-tests with 100,000 permutations to mitigate the risk of type I error. These $t$-tests compare the score distributions for each LLM (and GPT-4-Turbo) against GPT-3.5-Turbo across each coding preference and consistently reveal highly significant statistical differences (with $p<0.0001$ for most comparisons). Therefore, these results validate the discernible performance gaps between all models and GPT-3.5-Turbo. Moreover, the results also demonstrate a notably superior performance of GPT-4-Turbo relative to GPT-3.5-Turbo. This effectively illustrates GPT-3.5-Turbo’s impartiality as a judge, rewarding higher scores to more proficient LLMs and equitably evaluating its own responses.
In summary, our findings indicate that all LLMs, including highly capable ones such as WizardCoder-33B and DeepSeek-Coder-33B-Instruct, underperform compared to GPT-3.5-Turbo. This underperformance is likely due to a lack of alignment of the LLMs, and we believe fine-tuning them using alignment techniques might enhance their performance.
Figure 2: Win-tie-lose ratios of LLMs against GPT-3.5-Turbo. An LLM wins/loses if it gets a greater/lower score than GPT-3.5-Turbo. A tie is when the LLM and GPT-3.5-Turbo get identical scores.
III-B LLMs vs GPT-3.5-Turbo
@@ III-A LLMs Scores Exploration
IV Aligning LLMs to Coding Preferences
In this section, we present our experimental setup to improve LLM alignment with coding preferences using CodeUltraFeedback as preference data, which features the utilization of SFT and DPO [7, 38]. The purpose is twofold: validate the utility of CodeUltraFeedback for LLM alignment to coding preferences and show that a small LLM tuned using SFT and DPO can achieve greater alignment performance.
IV-A Supervised Fine-Tuning (SFT)
As demonstrated in Zephyr’s paper [38], an instruction-tuning phase is required prior to tuning an LLM using DPO as it facilitates tuning. This phase is achieved through supervised fine-tuning (SFT) using a dataset $\mathcal{D}=\{(x_{1},y_{1}),...,(x_{n},y_{n})\}$ of instruction-response pairs. Recent work leverages the Self-Instruct [46] and Evol-Instruct [43, 37] frameworks to compose such a dataset by generating instructions and responses using a highly capable model such as GPT-3.5/4 [33, 47, 40]. In this context, SFT is a form of knowledge distillation [48] that leverages a supervised signal from responses generated by a teacher LLM to tune a student LLM.
@@ IV-A Supervised Fine-Tuning (SFT)
IV-B Direct Preference Optimization (DPO)
@@ IV-C Experimental Details
Datasets. We build “Code Alignment Benchmark” (CODAL-Bench), a new benchmark for assessing LLM alignment to coding preferences. CODAL-Bench consists of 500 randomly selected samples from CodeUltraFeedback, with a balanced representation of each preference. The benchmark aims to provide a rigorous and comprehensive framework to evaluate and compare LLM alignment to coding preferences, allowing researchers to evaluate the impact of new alignment methods.
@@
For SFT, we use Magicoder Evol-Instruct [40], which consists of complex instruction-response pairs generated by GPT-4. We filter out the 10,000 samples used to build CodeUltraFeedback to mitigate data leakage between the SFT and DPO phases and avoid potential overfitting. The process results in 100,772 samples (95% for training, 5% for model evaluation).
@@
We use the remaining 9,500 samples from CodeUltraFeedback for DPO training. From each sample, we extract binary preferences ($y_{w}$, $y_{l}$), denoting the preferred and rejected response. These preferences are selected based on the highest and lowest ratings assigned by GPT-3.5-Turbo during the LLM-as-a-Judge annotation phase (see Section II-E). In instances where multiple responses have the highest/lowest ratings, we select one randomly to obtain $y_{w}$ / $y_{l}$.
Figure 3: Average alignment scores for LLMs across coding preferences on CODAL-Bench, evaluated using GPT-3.5-Turbo as a judge with reference-guided single-answer grading.
@@
LLMs. We conduct our experiments on CodeLlama-7B-Instruct as it comprises 7B parameters, allowing for tuning on a modest computing infrastructure. Furthermore, the model lags behind more proficient LLMs on CodeUltraFeedback (see Table III), which makes it an ideal candidate for demonstrating the potential impact of SFT and DPO on improving LLM alignment. For comparison, we include the following LLMs: CodeLlama-13/34B-Instruct, DeepSeek-Coder-6.7B-Instruct, WizardCoder-33B, GPT-3.5-Turbo, and GPT-4-Turbo.
@@
Grading Procedure. CODAL-Bench seeks to systematically compare LLM alignment, whereas CodeUltraFeedback’s main purpose is to serve as a training dataset for preference tuning. Therefore, we designed a new grading procedure adapted for evaluating each LLM individually using LLM-as-a-Judge. We implement a reference-guided single-answer grading system that has proven efficient in alignment benchmarks such as MT-Bench [34]. Each LLM’s response is individually evaluated against a reference response, providing a consistent basis for comparison and ensuring objective judgments across different LLMs. Additionally, we instruct the judge LLM to provide a rationale alongside the rating, enhancing the consistency of the evaluations. Ratings are assigned on a scale from 1 to 10, facilitating nuanced assessments of LLM alignment. We generate reference responses using GPT-3.5-Turbo and GPT-4-Turbo and use both models as judges.
@@
Training Details. We use HuggingFace TRL [49] to implement SFT and DPO with QLoRA [39] for efficient fine-tuning. We use a cosine learning rate scheduler with a learning rate of 2e-04 for SFT and 5e-05 for DPO, and 10% warmup steps. We fine-tune CodeLlama-7B-Instruct for 3 and 5 epochs for SFT and DPO, respectively. All experiments were conducted on a single NVIDIA RTX A5000 GPU (24GB). We provide extensive training details in our replication package.
V Results
Figure 4: Win-tie-lose ratios of LLMs against GPT-4-Turbo. DPO and SFT+DPO substantially increase the percentage of won and tie matches.
@@ V Results
GPT-4-Turbo stands out by surpassing all other LLMs across every coding preference by a considerable margin. Interestingly, all LLMs demonstrate lower scores on the instruction following preference. This trend suggests a relative challenge in achieving high performance in this area, possibly due to the inherently less precise nature of the preference and potentially vague instructions. Conversely, LLMs tend to achieve higher scores on code explanation and readability preferences We hypothesize this trend can be attributed to the fact that all LLMs have undergone instruction tuning, which can encourage LLMs to output reasoning and explanations alongside generated code. We observe that scores substantially increase when tuning CodeLlama-7B-Instruct with SFT+DPO, highlighting the effectiveness of these methods in enhancing model alignment across all coding preferences. For instance, the model yields a $7.89$ score on code complexity and efficiency preference, substantially surpassing CodeLlama-7B-Instruct ($6.16$) and larger models, including Wizard-33B ($6.93$) and CodeLlama-34B-Instruct ($6.42$).
@@
Different Judges and References Impact Alignment Scores. We investigate the effect of judges and references on the LLMs alignment scores in Table IV.
@@
SFT and DPO Improve Functional Correctness. While our findings highlight the positive impact of SFT and DPO on aligning LLMs with coding preferences, ensuring the functional correctness of the generated code remains a pivotal concern. In Table V, we report CodeLlama-7B-Instruct’s Pass@$k$ on HumanEval [1] and HumanEval+ [7].
Firstly, SFT substantially enhances Pass@$k$ on both benchmarks, with Pass@$1$ rising from 37.9 to 51.2 and 33.2 to 45.6 for HumanEval and HumanEval+, respectively. This result aligns with Magicoder [40] and WizardCoder [37] findings, showing that knowledge distillation through SFT substantially enhances the LLM’s effectiveness on these benchmarks. Secondly, while DPO and SFT+DPO variants exhibit a slight reduction in Pass@$k$ relative to SFT alone, they still maintain substantially higher Pass@$k$ compared to the baseline.
@@
TABLE IV: Average alignment scores of LLMs on CODAL-Bench. G-3.5 and G-4 refer to utilizing GPT-3.5-Turbo and GPT-4-Turbo responses as references, respectively.
@@
                         Reference:
G-3.5
G-4
G-4
CodeLlama-13B-Instruct
6.34
5.58
4.83
CodeLlama-34B-Instruct
6.49
5.84
5.36
DeepSeek-Coder-6.7B-Instruct
6.70
6.07
6.32
WizardCoder-33B
6.85
6.26
6.75
GPT-3.5-Turbo
7.18 7.35
GPT-4-Turbo 8.28
CodeLlama-7B-Instruct
6.00
5.46
4.72
   +SFT
6.51
5.83
5.84
   +DPO
7.15
6.79
5.08
@@
VI Discussion
In this paper, we have demonstrated how LLM-as-a-Judge enables the creation of new datasets and benchmarks like CodeUltraFeedback and CODAL-Bench to facilitate the tuning of LLMs for better alignment with coding preferences and assess their performance. This section offers further insights from our experiments, potential improvements for LLM alignment, and the significance of exploring varied LLM-as-a-Judge configurations. Additionally, we outline some limitations of CodeUltraFeedback and CODAL-Bench.
VI-A LLM Alignment and LLM-as-a-Judge
CodeUltraFeedback for Critic LLM Training. In this work, we demonstrate the utility of CodeUltraFeedback for LLM preference tuning using RLAIF and DPO. One of the drawbacks of CODAL-Bench is the need to leverage models like GPT-4 to evaluate other LLMs, which might turn out to be cost-prohibitive. Nonetheless, we believe the AI feedback data in CodeUltraFeedback, e.g., ratings and rationales, can be leveraged to fine-tune a small critic LLM trained to evaluate other LLMs. Prior work including Shepherd [50] and Prometheus [51] have been proposed around that idea and showed promising results.
TABLE V: Pass@$k$ of CodeLlama-7B-Instruct variants on HumanEval and HumanEval+.
@@ VI-A LLM Alignment and LLM-as-a-Judge
Parallel Findings with Zephyr Our findings draw parallels with Zephyr ’s [38] insights on DPO, with models beginning to overfit after a few epochs, achieving 100% training set accuracy. In our case, overfitting after five DPO training epochs correlated with improved performance. Although the model overfits, the gap between chosen and rejected rewards continued to grow, suggesting that monitoring reward accuracy might not be relevant for model selection. Therefore, more experimentation on DPO could illuminate better model tuning and selection approaches for improved model alignment.
Potential Judgment Biases. We highlight GPT-3.5-Turbo’s impartiality and robustness in judgment in both our dataset and benchmark. The performance ranking of LLMs remained consistent, underscoring the reliability of GPT-3.5-Turbo’s evaluations. Despite this, there remains the possibility of inherent biases in LLM judgments, such as a preference for lengthy or verbose responses. The grading procedures employed for CodeUltraFeedback and CODAL-Bench were carefully designed, drawing on established practices to preclude such biases [34, 52]. Single-answer grading with a reference response to the instruction and chain-of-thought prompting enables both GPT-3.5-Turbo and GPT-4-Turbo to produce consistent judgements. However, further exploration into the potential biases influencing LLM judgments is warranted.

### v2

@@ 3. CodeUltraFeedback: Dataset Analysis
In this section, we provide a detailed analysis of CodeUltraFeedback. Specifically, we explore the inherent capabilities of LLMs to align with coding preferences. Furthermore, we showcase the diversity of the dataset in terms of LLM responses and scores.
3.1. LLMs Scores Exploration
We start by exploring the ratings generated by GPT-3.5-Turbo in the LLM-as-a-Judge annotation phase. In Table 3, we present a detailed analysis of the average scores across coding preferences for each LLM, offering initial insight into their baseline capabilities.
At first glance, the scores might suggest marginal discrepancies given the 1–5 scoring range. However, these variations are substantial and carry statistical significance. To quantify this significance, we performed pairwise Welch’s $t$-tests, comparing the score distribution for each LLM against GPT-3.5-Turbo across each coding preference. The results reveal highly significant statistical differences, with $p<0.001$, for most comparisons and validate the discernible performance gaps between all models and GPT-3.5-Turbo. Moreover, the results also demonstrate a notably superior performance of GPT-4-Turbo relative to GPT-3.5-Turbo.
Overall, GPT-3.5-Turbo and GPT-4-Turbo score the highest across all preferences and on average. This outcome is expected considering both models have undergone extensive instruction and reinforcement learning from human feedback (RLHF) tunings to align more closely with human preferences (Ouyang et al., 2022). The scores indicate that highly capable LLMs such as WizardCoder-33B and DeepSeek-Coder-33B-Instruct underperform compared to GPT-3.5-Turbo. This underperformance is likely due to a lack of alignment of the LLMs, and tuning these models using alignment techniques could improve their performance.
3.2. LLMs Responses Analysis
Figure 2. Percentage of best and worst responses per LLM.
Figure 3. Distribution of best and worst responses length.
To gain more insights into the diversity of CodeUltraFeedback, we analyze the best and worst responses for each sample in terms of their assigned scores.
Fig. 2 depicts the percentage of best and worst responses per LLM in the dataset. Although GPT-4-Turbo and GPT-3.5-Turbo consistently score the highest across all coding preferences on average, this plot highlights that another LLM response is favoured for 73% of the samples. Moreover, even LLMs with lower average scores produce the highest-rated responses for at least 3.1% of the samples. This property of our dataset is valuable for downstream applications. For instance, it enables the alignment of LLMs using feedback data that encompasses a diverse range of models and mitigates the risk of the aligned LLM mimicking the behaviour of a single model.
Next, we analyze the distribution of best and worst response lengths in Fig. 3. Both distributions show similar patterns, though the best responses tend to be slightly longer. Importantly, both distributions encompass a broad spectrum of response lengths. This characteristic of CodeUltraFeedback is also valuable for model alignment, as it helps prevent the aligned LLM from developing a bias towards excessive verbosity.
3.3. LLMs vs GPT-3.5-Turbo
@@ 3.2. LLMs Responses Analysis
Figure 4. Win-tie-lose ratios of LLMs against GPT-3.5-Turbo. An LLM wins/loses if it gets a greater/lower score than GPT-3.5-Turbo. A tie is when the LLM and GPT-3.5-Turbo get identical scores.
4. Aligning LLMs to Coding Preferences
In this section, we present our experimental setup to improve LLM alignment with coding preferences using CodeUltraFeedback as preference data, which features the utilization of SFT and DPO (Liu et al., 2023b; Tunstall et al., 2023). The purpose is twofold: validate the utility of CodeUltraFeedback for LLM alignment to coding preferences and show that a small LLM tuned using SFT and DPO can achieve greater alignment performance.
4.1. Supervised Fine-Tuning (SFT)
As demonstrated in Zephyr’s paper (Tunstall et al., 2023), an instruction-tuning phase is required prior to tuning an LLM using DPO as it facilitates tuning. This phase is achieved through supervised fine-tuning (SFT) using a dataset $\mathcal{D}=\{(x_{1},y_{1}),...,(x_{n},y_{n})\}$ of instruction-response pairs. Recent work leverages the Self-Instruct (Wang et al., 2022a) and Evol-Instruct (Xu et al., 2023; Luo et al., 2023) frameworks to compose such a dataset by generating instructions and responses using a highly capable model such as GPT-3.5/4 (Ouyang et al., 2022; Taori et al., 2023; Wei et al., 2023). In this context, SFT is a form of knowledge distillation (Xu et al., 2024) that leverages a supervised signal from responses generated by a teacher LLM to tune a student LLM.
@@ 4.1. Supervised Fine-Tuning (SFT)
4.2. Direct Preference Optimization (DPO)
@@ 4.3. Experimental Details
Datasets. We build CodeUltraFeedback-Bench, a new benchmark for assessing LLM alignment to coding preferences. CodeUltraFeedback-Bench consists of 500 randomly selected samples from CodeUltraFeedback, with a balanced representation of each preference. The benchmark aims to provide a rigorous and comprehensive framework to evaluate and compare LLM alignment to coding preferences, allowing researchers to evaluate the impact of new alignment methods.
@@
For SFT, we use Magicoder Evol-Instruct (Wei et al., 2023), which consists of instruction-response pairs generated by GPT-4. We filter out the 10,000 samples used to build CodeUltraFeedback to mitigate data leakage between the SFT and DPO phases and avoid potential overfitting. The process results in 100,772 samples (95% for training, 5% for model evaluation).
@@
We use the remaining 9,500 samples from CodeUltraFeedback for DPO. From each sample, we extract binary preferences ($y_{w}$, $y_{l}$), denoting the preferred and rejected response. These preferences are selected based on the highest and lowest ratings assigned by GPT-3.5-Turbo during the LLM-as-a-Judge annotation phase (see Section 2.5). In instances where multiple responses have the highest/lowest ratings, we select one randomly to obtain $y_{w}$ / $y_{l}$.
@@
LLMs. We conduct our experiments on CodeLlama-7B-Instruct as it comprises 7B parameters, allowing for tuning on a modest computing infrastructure. Furthermore, the model lags behind more proficient LLMs in CodeUltraFeedback’s scores (see Table 3), which makes it an ideal candidate for demonstrating the potential impact of SFT and DPO on improving LLM alignment. For comparison, we include the following LLMs: CodeLlama-13/34B-Instruct, DeepSeek-Coder-6.7B-Instruct, WizardCoder-33B, GPT-3.5-Turbo, and GPT-4-Turbo.
@@
Grading Procedure. CodeUltraFeedback-Bench seeks to systematically compare LLM alignment, whereas CodeUltraFeedback’s main purpose is to serve as a training dataset for preference tuning. Therefore, we designed a new grading procedure adapted for evaluating each LLM individually using LLM-as-a-Judge. We implement a reference-guided single-answer grading system that has proven efficient in alignment benchmarks such as MT-Bench (Zheng et al., 2023a). Each LLM’s response is individually evaluated against a reference response, providing a consistent basis for comparison and ensuring objective judgments across LLMs. Additionally, we instruct the judge LLM to provide a rationale alongside the rating, enhancing the consistency of the evaluations. Ratings are assigned on a scale from 1 to 10, facilitating nuanced assessments of LLM alignment. We generate reference responses using GPT-3.5-Turbo and GPT-4-Turbo and use both models as judges.
@@
Training Details. We use HuggingFace TRL (Leandro et al., 2020) to implement SFT and DPO with QLoRA (Dettmers et al., 2023) for efficient fine-tuning. We use a cosine learning rate scheduler with a learning rate of 2e-04 for SFT and 5e-05 for DPO, and 10% warmup steps. We fine-tune CodeLlama-7B-Instruct for 3 and 5 epochs for SFT and DPO, respectively. All experiments were conducted on a single NVIDIA RTX A5000 GPU (24GB). We provide extensive training details in our replication package.
5. Results
Figure 5. Average alignment scores for LLMs across coding preferences on CodeUltraFeedback-Bench, evaluated using GPT-3.5-Turbo as a judge with reference-guided single-answer grading.
@@ 5. Results
GPT-4-Turbo stands out by surpassing all other LLMs across every coding preference by a considerable margin. Interestingly, all LLMs demonstrate lower scores on the instruction following preference. This trend suggests a relative challenge in achieving high performance in this area, possibly due to the inherently less precise nature of the preference and potentially vague instructions. Conversely, LLMs tend to achieve higher scores on code explanation and readability preferences. We hypothesize this trend can be attributed to the fact that all LLMs have undergone instruction tuning, which can encourage LLMs to output reasoning and explanations alongside generated code. We observe that scores substantially increase when tuning CodeLlama-7B-Instruct with SFT+DPO, highlighting the effectiveness of these methods in enhancing model alignment across all coding preferences. For instance, the model yields a $7.89$ score on code complexity and efficiency preference, substantially surpassing CodeLlama-7B-Instruct ($6.16$) and larger models, including Wizard-33B ($6.93$) and CodeLlama-34B-Instruct ($6.42$).
@@
Figure 6. Win-tie-lose ratios of LLMs against GPT-4-Turbo. DPO and SFT+DPO substantially increase the percentage of won and tie matches.
@@
Different Judges and References Impact Alignment Scores. We investigate the effect of judges and references on the LLMs alignment scores in Table 4.
@@
SFT and DPO Improve Functional Correctness. While our findings highlight the positive impact of SFT and DPO on aligning LLMs with coding preferences, ensuring the functional correctness of the generated code remains a pivotal concern. In Table 5, we report CodeLlama-7B-Instruct’s Pass@$k$ on HumanEval (Chen and others., 2021) and HumanEval+ (Liu et al., 2023b).
Firstly, SFT substantially enhances Pass@$k$ on both benchmarks, with Pass@$1$ rising from 37.9 to 51.2 and 33.2 to 45.6 for HumanEval and HumanEval+, respectively. This result aligns with Magicoder (Wei et al., 2023) and WizardCoder (Luo et al., 2023) findings, showing that knowledge distillation through SFT substantially enhances the LLM’s effectiveness on these benchmarks. Secondly, while DPO and SFT+DPO variants exhibit a slight reduction in Pass@$k$ relative to SFT alone, they still maintain substantially higher Pass@$k$ compared to the baseline.
@@
Table 4. Average alignment scores of LLMs on CodeUltraFeedback-Bench. G-3.5 and G-4 refer to utilizing GPT-3.5-Turbo and GPT-4-Turbo responses as references, respectively.
@@
                          Reference G-3.5 G-4 G-4
CodeLlama-13B-Instruct 6.34 5.58 4.83
CodeLlama-34B-Instruct 6.49 5.84 5.36
DeepSeek-Coder-6.7B-Instruct 6.70 6.07 6.32
WizardCoder-33B 6.85 6.26 6.75
GPT-3.5-Turbo - 7.18 7.35
GPT-4-Turbo 8.28 - -
CodeLlama-7B-Instruct 6.00 5.46 4.72
+SFT 6.51 5.83 5.84
+DPO 7.15 6.79 5.08
@@
Table 5. Pass@$k$ of CodeLlama-7B-Instruct variants on HumanEval and HumanEval+.
@@
6. Discussion
In this paper, we have demonstrated how LLM-as-a-Judge enables the creation of new datasets and benchmarks like CodeUltraFeedback and CodeUltraFeedback-Bench to facilitate the tuning of LLMs for better alignment with coding preferences and assess their performance. This section offers further insights from our experiments, potential improvements for LLM alignment, and the significance of exploring varied LLM-as-a-Judge configurations. Additionally, we outline some limitations of CodeUltraFeedback and CodeUltraFeedback-Bench.
6.1. LLM Alignment and LLM-as-a-Judge
CodeUltraFeedback for Critic LLM Training. In this work, we demonstrate the utility of CodeUltraFeedback for LLM preference tuning using RLAIF and DPO. One of the drawbacks of CodeUltraFeedback-Bench is the need to leverage models like GPT-4 to evaluate other LLMs, which might turn out to be cost-prohibitive. Nonetheless, we believe the AI feedback data in CodeUltraFeedback, e.g., ratings and rationales, can be leveraged to fine-tune a small critic LLM trained to evaluate other LLMs. Prior work including Shepherd (Wang et al., 2023) and Prometheus (Kim et al., 2023) have been proposed around that idea and showed promising results.
Parallel Findings with Zephyr Our findings draw parallels with Zephyr ’s (Tunstall et al., 2023) insights on DPO, with models beginning to overfit after a few epochs, achieving 100% training set accuracy. In our case, overfitting after five DPO training epochs correlated with improved performance. Although the model overfits, the gap between chosen and rejected rewards continued to grow, suggesting that monitoring reward accuracy might not be relevant for model selection. Therefore, more experimentation on DPO could illuminate better model tuning and selection approaches for improved model alignment.
Potential Judgment Biases. We highlight GPT-3.5-Turbo’s robustness in judgment in both our dataset and benchmark. The performance ranking of LLMs remained consistent, underscoring the reliability of GPT-3.5-Turbo’s evaluations. Despite this, there remains the possibility of inherent biases in LLM judgments, such as a preference for lengthy or verbose responses. The grading procedures employed for CodeUltraFeedback and CodeUltraFeedback-Bench were carefully designed, drawing on established practices to preclude such biases (Zheng et al., 2023a; Li et al., 2023b). Single-answer grading with a reference response to the instruction and chain-of-thought prompting enables both GPT-3.5-Turbo and GPT-4-Turbo to produce consistent judgements. However, further exploration into the potential biases influencing LLM judgments is warranted.